OpenStudy (anonymous):

e^3*ln(x^2)*2(ln)(x) =ln (e^16)

OpenStudy (anonymous):

howdy!

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

see that ln (x^2) ? that can be written in terms of just ln x.... and the right side? simplify that.

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

yes definitely Im just learning this

OpenStudy (anonymous):

ok... the ln (x^2) can be written as 2ln(x).. it's the property: ln(a^b) = bln(a)

OpenStudy (anonymous):

and that rigt side... ln(e^16) = ....

OpenStudy (anonymous):

16 ln e

OpenStudy (anonymous):

right....

OpenStudy (anonymous):

do you know what ln e is?

OpenStudy (anonymous):

the number its associated with? 2.71..?

OpenStudy (anonymous):

well e=2.71.... but what about ln(e) ?

OpenStudy (anonymous):

they are equal to each other?

OpenStudy (anonymous):

no, ln e = 1.

OpenStudy (anonymous):

remember the base of ln is e correct?

OpenStudy (anonymous):

it's like asking the question:  what's x?